# Weighted Orlicz *-algebras on locally elliptic groups

### v1

@@ 1. Introduction
In the case of connected Lie groups, it is well known that every connected nilpotent Lie group is Hermitian, Wiener and Boidol, and for connected solvable Lie groups of real dimension $\leqslant 4$, there is one exceptional group, the “Poguntke group” [50, §$\S$12.6.27], which does not satisfy these properties. On the other hand, non-compact connected semisimple Lie groups are never Hermitian, Wiener or Boidol [50, Chapter 12].
@@
To be more explicit, one constructs Orlicz spaces as follows: given a measure space $(X,\mu)$ and a Young function $\Phi:\mathbb{R}\rightarrow[0,\infty]$, one can associate a certain Banach space $L^{\Phi}(X)$ of measurable complex valued functions on $X$, and it is this space that we call an Orlicz space (see Section 2.2 for the complete definition or [55] for further theoretical details). The case when $\Phi(x)=\lvert x\rvert^{p}/p$ gives the classical $L^{p}$-spaces. Now, given a locally compact group $G$ and a Young function $\Phi$, one can study the space $L^{\Phi}(G)$. Typically $L^{\Phi}(G)$ is not closed under convolution, even when this is an $L^{p}$-space, but for a suitable weight $\omega$ on $G$, the weighted Orlicz space $L^{\Phi}(G,\omega)$ may be a Banach *-algebra when equipped with the convolution product and the involution $f^{*}(x):=\overline{f(x^{-1})}$ ($f\in L^{\Phi}(G,\omega)$). This is what we will refer to as a weighted Orlicz *-algebra (or sometimes a weighted $L^{\Phi}$-algebra) and we will call such a weight $\omega$ an $L^{\Phi}$-weight.
@@
This result could be useful in the study of the unitary representation theory of locally elliptic groups. Indeed, for example, it is currently an open question for which locally elliptic groups $G$ is the topology on $\text{\rm Prim}(C^{*}(G))$ $T_{1}$. Sometimes $C^{*}(G)$ isn’t the most convenient algebra to work with: being a completion of $L^{1}(G)$, one cannot treat all elements of $C^{*}(G)$ as functions on $G$. Thus, it is often convenient to work with $L^{1}(G)$ instead. In the case of a locally elliptic group, as given by the above theorem, $\text{\rm Prim}_{*}(L^{1}(G))$ and $\text{\rm Prim}(C^{*}(G))$ are homeomorphic, so it is fine working with $L^{1}(G)$ instead of $C^{*}(G)$ in the study of the unitary representation theory. However, unlike $C^{*}(G)$, $L^{1}(G)$ may not be Hermitian, which is a desirable property to have. But even if $L^{1}(G)$ is not Hermitian, as is shown in the article, there always exist weights on $G$ such that $L^{1}(G,\omega)$ is Hermitian (and $\text{\rm Prim}_{*}(L^{1}(G,\omega))\cong\text{\rm Prim}(C^{*}(G)))$. So in the context of non-Hermitian locally elliptic groups, working with a weighted $L^{1}$-algebra may be more natural algebra than working with $C^{*}(G)$ or $L^{1}(G)$ when studying the unitary representation theory of these groups.
@@ 2.2. Orlicz spaces
[…]
@@
[…]
We note that in the case that $\Phi(x)=\lvert x\rvert^{p}/p$ ($p\in(1,\infty)$), the complementary Young function to $\Phi$ is $\Psi=\lvert x\rvert^{q}/q$ where $11/p+1/q=$ 1, and these Young functions satisfy the $\Delta_{2}$ condition. Furthermore, in this case, $L^{\Phi}(G)$ becomes the classical $L^{p}$-space denoted $L^{p}(G)$, and the Orlicz norm with respect to the Young function $\Phi$ is equivalent to the usual $L^{p}$-norm on $L^{p}(G)$.
@@
Let $\Phi$ be a Young function and $\omega$ a weight on a locally compact group $G$. We call $\omega$ an $L^{\Phi}$-weight if $L^{\Phi}(G,\omega)$ is a Banach *-algebra when equipped with the convolution product and the involution $f^{*}(x)=\Delta(x)\overline{f(x^{-1})}$, where $\Delta$ is the modular function on $G$.
Remark that, by our definition of a weight function, if $G$ is a unimodular group then the space $L^{1}(G,\omega)$ is always a Banach *-algebra. In particular, on a unimodular locally compact group, every weight is an $L^{1}$-weight.
@@
[47, 58] Let $G$ be a locally compact group with left Haar measure $\mu$, $(\Phi,\Psi)$ a complementary pair of Young functions and $\omega$ a $L^{\Phi}$-weight on $G$. The following hold:
@@ 2.3. Locally elliptic locally compact groups
Of course, if $G$ is discrete, then locally elliptic is synonymous with the term locally finite.
@@ 3. Weights on locally elliptic groups
[14, Korollar 3.8] Let $G$ be a locally compact unimodular group and $\omega$ a locally-integrable sub-additive weight on $G$. If $p\in(1,\infty)$ and $\omega^{-1}\in L^{q}(G)$ where $q:=\frac{p}{p-1}$, then $L^{p}(G,\omega)$ is a Banach *-algebra.
@@
[…]
@@
For $n\in\mathbb{N}$, let $C_{n}$ denote the cyclic group of order $n$. Define the group $G:=\bigoplus_{n=1}^{\infty}C_{n}$. Clearly $G$ is a locally finite group and hence locally elliptic. We define a weight $\omega$ on $G$ as follows: let $x_{n}$ denote the canonical generator of the copy of $C_{n}$ in $G$ for each $n\in\mathbb{N}$. Then, given a general element $x=x_{n_{1}}^{m_{1}}\cdots x_{n_{k}}^{m_{k}}\in G$, where $11\leqslant m_{i}\leqslant n_{i}-$ 1 is an integer for each $i$, we define $\omega$ on $x$ by
@@ 4. The Hermitian property
Also, it would be interesting to have characterisations of Hermitianness for $L^{1}(G,\omega)$ which are independent of the Hermitian property of $L^{1}(G)$, when $G$ is a locally elliptic group. Although we can’t say anything in general at the moment, we note the following result and an immediate corollary of it.
@@
[…]
@@
Let $G$ be a locally elliptic group and $\omega$ a weight on $G$ that satisfies the uniform GRS condition. Then, for all $f\in L^{1}(G)$, $\sigma_{L^{1}(G,\omega)}(f)=\sigma_{L^{1}(G)}(f)$.
@@
[…]
@@
[…]
@@ 5. Functional calculus and the Wiener property
Given $10<\gamma<$ 1, define a weight on $\mathbb{Z}$ by $\omega_{\gamma}(n):=e^{2\lvert n\rvert^{\gamma}}$ ($n\in\mathbb{Z}$). Let $\tilde{A}_{\gamma}$ denote the algebra of $2\pi$-periodic $C^{\infty}$-functions $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ with Fourier coefficients in $\ell^{1}(\mathbb{Z},\omega_{\gamma})$ under pointwise multiplication. In particular, the algebra $\tilde{A}_{\gamma}$ consists of functions of the form $\varphi(x)=\sum_{n\in\mathbb{Z}}\hat{\varphi}(n)e^{inx}$ such that $\sum_{n\in\mathbb{Z}}\lvert\hat{\varphi}(n)\rvert e^{2\lvert n\rvert^{\gamma}}<\infty$. Then, we let $A_{\gamma}$ be the subalgebra of $\tilde{A}_{\gamma}$ consisting of those functions $\varphi\in\tilde{A}_{\gamma}$ with $\varphi(0)=0$.
@@
[12] For $10<\gamma<$ 1, the algebra $A_{\gamma}$ contains functions of arbitrarily small support. Furthermore, for every $\epsilon>0$ and every interval $[p,q]\subseteq(0,2\pi)$ with $p+\epsilon<q-\epsilon$, there exists a function $\varphi\in A_{\gamma}$ satisfying the following properties:
@@ 6. The *-regular property and representation theory of locally elliptic groups
$[-\sup_{\pi\in C}\left\lVert\pi(g)\right\rVert_{\mathcal{B}(\mathcal{H}_{\pi})},\sup_{\pi\in C}\left\lVert\pi(g)\right\rVert_{\mathcal{B}(\mathcal{H}_{\pi})}]\subseteq[-1,1]$
@@
    $\text{\rm Prim}(L^{\Phi}(G,\omega))=\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))=\text{\rm Max}(L^{\Phi}(G,\omega))$.
@@
The equivalence of $(i)$, $(ii)$ and $(iii)$ follows directly from Theorem 1.5. Clearly $(v)$ implies $(iv)$. We now show that $(iv)$ implies $(v)$. If $M\in\text{\rm Max}(L^{\Phi}(G,\omega))$, then $L^{\Phi}(G,\omega)/M$ is a simple $L^{\Phi}(G,\omega)$-module which has annihilator $M$, thus $\text{\rm Max}(L^{\Phi}(G,\omega))\subseteq\text{\rm Prim}(L^{\Phi}(G,\omega))$. Since $L^{\Phi}(G,\omega)$ is Hermitian, $\text{\rm Prim}(L^{\Phi}(G,\omega))\subseteq\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$, and since $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))\subseteq\text{\rm Max}(L^{\Phi}(G,\omega))$ by the assumption of $(iv)$, $(v)$ follows.
Finally, we need to prove the equivalence of $(iii)$ and $(iv)$. First, let’s suppose that $(iii)$ is true and prove $(iv)$. Suppose for a contradiction that there is an ideal $J$ in $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$ which is not maximal. Then, there exists a proper closed (two-sided) ideal $J^{\prime}\subseteq L^{\Phi}(G,\omega)$ properly containing $J$. Since $L^{\Phi}(G,\omega)$ is Hermitian and hence Wiener by Theorem 1.3, $\text{\rm hull}_{*}(\{J^{\prime}\})$ is non-empty and hence contains an ideal $I$ in $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$. In particular, $\text{\rm hull}_{*}(\{J\})\supseteq\{I,J\}$, from which it follows, by definition of the hull-kernel topology, that $J$ is not a closed point of $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$. This contradicts the fact that $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$ is $T_{1}$.
@@
It is an ongoing area of research to determine which locally elliptic groups $G$ is $\text{\rm Prim}(C^{*}(G))$ $T_{1}$. We note that if $G$ is either 2-step-nilpotent, or nilpotent and has a compact open normal subgroup, then it is already known that $\text{\rm Prim}(C^{*}(G))$ is $T_{1}$ [52, 7].
@@ 7. Minimal ideals of a given hull
[…]
@@
$[-\sup_{\text{\rm ker}(\pi)\in C}\left\lVert\pi(f)\right\rVert_{B(\mathcal{H}_{\pi})},\sup_{\text{\rm ker}(\pi)\in C}\left\lVert\pi(f)\right\rVert_{B(\mathcal{H}_{\pi})}]$

### v2

@@ 1. Introduction
In the case of connected Lie groups, it is well known that every connected nilpotent Lie group is Hermitian, Wiener and Boidol, and for connected solvable Lie groups of real dimension $\leqslant 4$, there is one exceptional group, the “Poguntke group” [50, $\S$12.6.27], which does not satisfy these properties. On the other hand, non-compact connected semisimple Lie groups are never Hermitian, Wiener or Boidol [50, Chapter 12].
@@
To be more explicit, one constructs Orlicz spaces as follows: given a measure space $(X,\mu)$ and a Young function $\Phi:\mathbb{R}\rightarrow[0,\infty]$, one can associate a certain Banach space $L^{\Phi}(X)$ of measurable complex valued functions on $X$, and it is this space that we call an Orlicz space (see Section 2.2 for the complete definition or [55] for further theoretical details). The case when $\Phi(x)=\lvert x\rvert^{p}/p$ gives the classical $L^{p}$-spaces. Now, given a locally compact group $G$ and a Young function $\Phi$, one can study the space $L^{\Phi}(G)$. Typically $L^{\Phi}(G)$ is not closed under convolution, even when this is an $L^{p}$-space, but if $G$ is unimodular and $\omega$ a suitable weight on $G$, the weighted Orlicz space $L^{\Phi}(G,\omega)$ may be a Banach *-algebra when equipped with the convolution product and the involution $f^{*}(x):=\overline{f(x^{-1})}$ ($f\in L^{\Phi}(G,\omega)$). This is what we will refer to as a weighted Orlicz *-algebra (or sometimes a weighted $L^{\Phi}$-algebra) and we will call such a weight $\omega$ an $L^{\Phi}$-weight.
@@
This result could be useful in the study of the unitary representation theory of locally elliptic groups. Indeed, for example, it is currently an open question for which locally elliptic groups $G$ does the topology on $\text{\rm Prim}(C^{*}(G))$ satisfy the $T_{1}$ separation axiom. Sometimes $C^{*}(G)$ is not the most convenient algebra to work with: being a completion of $L^{1}(G)$, one cannot treat all elements of $C^{*}(G)$ as functions on $G$. Thus, it is often convenient to work with $L^{1}(G)$ instead. In the case of a locally elliptic group, as given by the above theorem, $\text{\rm Prim}_{*}(L^{1}(G))$ and $\text{\rm Prim}(C^{*}(G))$ are homeomorphic, so it is fine working with $L^{1}(G)$ instead of $C^{*}(G)$ in the study of the unitary representation theory. However, unlike $C^{*}(G)$, $L^{1}(G)$ may not be Hermitian, which is a desirable property to have. But even if $L^{1}(G)$ is not Hermitian, as is shown in the article, there always exist weights on $G$ such that $L^{1}(G,\omega)$ is Hermitian (and $\text{\rm Prim}_{*}(L^{1}(G,\omega))\cong\text{\rm Prim}(C^{*}(G)))$. So in the context of non-Hermitian locally elliptic groups, working with a weighted $L^{1}$-algebra may be more natural algebra than working with $C^{*}(G)$ or $L^{1}(G)$ when studying the unitary representation theory of these groups.
@@ 2.2. Orlicz spaces
$L^{\Phi}(G):=\bigg\{f:G\rightarrow\mathbb{C}\text{ measurable}:\int_{G}\Phi(\alpha\lvert f\rvert)\>dx<\infty\text{ for some }\alpha>0\bigg\}.$
@@
$N_{\Phi}(f):=\inf\bigg\{k>0:\int_{G}\Phi(\lvert f(x)\rvert/k)\>dx\leqslant 1\bigg\}\;\;\;\;(f\in L^{\Phi}(G)).$
We note that in the case that $\Phi(x)=\lvert x\rvert^{p}/p$ ($p\in(1,\infty)$), the complementary Young function to $\Phi$ is $\Psi=\lvert x\rvert^{q}/q$ where $1/p+1/q=1$, and these Young functions satisfy the $\Delta_{2}$ condition. Furthermore, in this case, $L^{\Phi}(G)$ becomes the classical $L^{p}$-space denoted $L^{p}(G)$, and the Orlicz norm with respect to the Young function $\Phi$ is equivalent to the usual $L^{p}$-norm on $L^{p}(G)$.
@@
Let $\Phi$ be a Young function and $\omega$ a weight on a unimodular locally compact group $G$. We call $\omega$ an $L^{\Phi}$-weight if $L^{\Phi}(G,\omega)$ is a Banach *-algebra when equipped with the convolution product and the involution $f^{*}(x)=\overline{f(x^{-1})}$.
@@
[47, 58] Let $G$ be a unimodular locally compact group with left Haar measure $\mu$, $(\Phi,\Psi)$ a complementary pair of Young functions and $\omega$ a $L^{\Phi}$-weight on $G$. The following hold:
@@ 2.3. Locally elliptic locally compact groups
Of course, if $G$ is discrete, then locally elliptic is synonymous with the term locally finite. One also checks easily that every locally elliptic group is unimodular.
@@ 3. Weights on locally elliptic groups
[14, Korollar 3.8] Let $G$ be a unimodular locally compact group and $\omega$ a locally-integrable sub-additive weight on $G$. If $p\in(1,\infty)$ and $\omega^{-1}\in L^{q}(G)$ where $q:=\frac{p}{p-1}$, then $L^{p}(G,\omega)$ is a Banach *-algebra.
@@
$\int_{G}(1/\omega^{\sharp}_{p}(x))^{q}\>d\mu(x)$ $=\int_{K_{1}}(1/\omega^{\sharp}_{p}(x))^{q}\>d\mu(x)+\sum_{i=2}^{\infty}\int_{K_{i}\smallsetminus K_{i-1}}(1/\omega^{\sharp}_{p}(x))^{q}\>d\mu(x)$
@@
For $n\in\mathbb{N}$, let $C_{n}$ denote the cyclic group of order $n$. Define the group $G:=\bigoplus_{n=1}^{\infty}C_{n}$. Clearly $G$ is a locally finite group and hence locally elliptic. We define a weight $\omega$ on $G$ as follows: let $x_{n}$ denote the canonical generator of the copy of $C_{n}$ in $G$ for each $n\in\mathbb{N}$. Then, given a general element $x=x_{n_{1}}^{m_{1}}\cdots x_{n_{k}}^{m_{k}}\in G$, where $1\leqslant m_{i}\leqslant n_{i}-1$ is an integer for each $i$, we define $\omega$ on $x$ by
@@ 4. The Hermitian property
Also, it would be interesting to have characterisations of Hermitianness for $L^{1}(G,\omega)$ which are independent of the Hermitian property of $L^{1}(G)$, when $G$ is a locally elliptic group. Although we cannot say anything in general at the moment, we note the following result and an immediate corollary of it.
@@
$\left\lVert\omega_{\bf{a}}^{-1}\right\rVert_{L^{p}(G)}^{p}=\int_{G}\lvert\omega_{\bf{a}}^{-1}(x)\rvert^{p}\>d\mu(x)$
$=\int_{K_{1}}\lvert\omega_{\bf{a}}^{-1}(x)\rvert^{p}\>d\mu(x)+\sum_{i=2}^{\infty}\int_{K_{i}\smallsetminus K_{i-1}}\lvert\omega_{\bf{a}}^{-1}(x)\rvert^{p}\>d\mu(x)$
@@
Let $G$ be a locally elliptic group and $\omega$ a weight on $G$ that satisfies the uniform GRS condition. Then, for all $f\in L^{1}(G,\omega)$, $\sigma_{L^{1}(G,\omega)}(f)=\sigma_{L^{1}(G)}(f)$.
@@
$\left\lVert f^{*n}\right\rVert_{L^{1}(G,\omega)}\leqslant\int_{G}\cdots\int_{G}\lvert f(x_{1})\rvert\cdots\lvert f(x_{n})\rvert\omega(x_{1}x_{2}\cdots x_{n})dx_{1}dx_{2}\cdots dx_{n}$
$=\sum_{i_{1},\dots,i_{n}=0}^{\infty}\int_{K_{i_{1}}\smallsetminus K_{i_{1}-1}}\cdots\int_{K_{i_{n}}\smallsetminus K_{i_{n}-1}}\lvert f(x_{1})\rvert\cdots\lvert f(x_{n})\rvert\omega(x_{1}x_{2}\cdots x_{n})dx_{1}dx_{2}\cdots dx_{n}$
@@
Then, let $a_{i}:=\int_{K_{i}\smallsetminus K_{i-1}}\lvert f(x)\rvert\>dx$ for each $i\in\mathbb{N}\cup\{0\}$ and $a:=(a_{n})_{n\in\mathbb{N}\cup\{0\}}$. Clearly $\left\lVert f\right\rVert_{L^{1}(G)}=\left\lVert a\right\rVert_{\ell^{1}(\mathbb{Z})}$. It then follows by the previous arguments that
@@ 5. Functional calculus and the Wiener property
Given $0<\gamma<1$, define a weight on $\mathbb{Z}$ by $\omega_{\gamma}(n):=e^{2\lvert n\rvert^{\gamma}}$ ($n\in\mathbb{Z}$). Let $\tilde{A}_{\gamma}$ denote the algebra of $2\pi$-periodic $C^{\infty}$-functions $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ with Fourier coefficients in $\ell^{1}(\mathbb{Z},\omega_{\gamma})$ under pointwise multiplication. In particular, the algebra $\tilde{A}_{\gamma}$ consists of functions of the form $\varphi(x)=\sum_{n\in\mathbb{Z}}\hat{\varphi}(n)e^{inx}$ such that $\sum_{n\in\mathbb{Z}}\lvert\hat{\varphi}(n)\rvert e^{2\lvert n\rvert^{\gamma}}<\infty$. Then, we let $A_{\gamma}$ be the subalgebra of $\tilde{A}_{\gamma}$ consisting of those functions $\varphi\in\tilde{A}_{\gamma}$ with $\varphi(0)=0$.
@@
[12] For $0<\gamma<1$, the algebra $A_{\gamma}$ contains functions of arbitrarily small support. Furthermore, for every $\epsilon>0$ and every interval $p,q]\subseteq(0,2\pi)$ with $p+\epsilon<q-\epsilon$, there exists a function $\varphi\in A_{\gamma}$ satisfying the following properties:
@@ 6. The *-regular property and representation theory of locally elliptic groups
$\sup_{\pi\in C}\left\lVert\pi(g)\right\rVert_{\mathcal{B}(\mathcal{H}_{\pi})},\sup_{\pi\in C}\left\lVert\pi(g)\right\rVert_{\mathcal{B}(\mathcal{H}_{\pi})}]\subseteq[-1,1]$
@@
    $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))=\text{\rm Max}(L^{\Phi}(G,\omega))$.
@@
The equivalence of $(i)$, $(ii)$ and $(iii)$ follows directly from Theorem 1.5. Clearly $(v)$ implies $(iv)$. Since $L^{\Phi}(G,\omega)$ is Hermitian by assumption, it is Wiener by Theorem 1.3, which implies that $\text{\rm Max}(L^{\Phi}(G,\omega))\subseteq\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$. Thus $(iv)$ implies $(v)$.
Finally, we need to prove the equivalence of $(iii)$ and $(iv)$. First, let’s suppose that $(iii)$ is true and prove $(iv)$. Suppose for a contradiction that there is an ideal $J$ in $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$ which is not maximal amongst closed two-sided ideals. Then, there exists a proper closed (two-sided) ideal $J^{\prime}\subseteq L^{\Phi}(G,\omega)$ properly containing $J$. Since $L^{\Phi}(G,\omega)$ is Hermitian and hence Wiener by Theorem 1.3, $\text{\rm hull}_{*}(\{J^{\prime}\})$ is non-empty and hence contains an ideal $I$ in $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$. In particular, $\text{\rm hull}_{*}(\{J\})\supseteq\{I,J\}$, from which it follows, by definition of the hull-kernel topology, that $J$ is not a closed point of $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$. This contradicts the fact that $\text{\rm Prim}_{*}(L^{\Phi}(G,\omega))$ is $T_{1}$.
@@
It is an ongoing area of research to determine for which locally elliptic groups $G$ does the topology on $\text{\rm Prim}(C^{*}(G))$ satisfy the $T_{1}$ separation axiom. We note that if $G$ is either 2-step-nilpotent, or nilpotent and has a compact open normal subgroup, then it is already known that $\text{\rm Prim}(C^{*}(G))$ is $T_{1}$ [52, 7].
@@ 7. Minimal ideals of a given hull
$m(C):=\{$ $\varphi\{f\}:f=f^{*}\in C_{c}(G),\left\lVert f\right\rVert_{L^{1}(G)}\leqslant 1,\varphi\in A_{\gamma},\varphi=0\text{ on a}$
@@
$\sup_{\text{\rm ker}(\pi)\in C}\left\lVert\pi(f)\right\rVert_{B(\mathcal{H}_{\pi})},\sup_{\text{\rm ker}(\pi)\in C}\left\lVert\pi(f)\right\rVert_{B(\mathcal{H}_{\pi})}]$